OpenStudy (anonymous):

Steps to w^2-2w=24+3w

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

First of all your question "Steps to" doesn't mean anything?

OpenStudy (skullpatrol):

Are you looking for the steps to solving $$w^2-2w=24+3w$$?

OpenStudy (skullpatrol):

To solve an equation you are looking for the solution set of the equation. That is the set of all solutions of the sentence that make it a true statement.

OpenStudy (skullpatrol):

In this case we will use the zero-product property: For all real numbers a and b: $$ab=0$$ if and only if a=0 or b=0. A product of factors is zero if and only if one or more of the factors is zero.

OpenStudy (skullpatrol):

To use the zero-product property we must put your equation: $$ w^2-2w=24+3w$$ into the form: $$ab=0$$ and then find what values will make it a true statement. That is the set of all solutions of the sentence that make it a true statement.

OpenStudy (skullpatrol):

The first step is to bring all the terms to one side and thus making the other side equal to 0 just like ab=0 requires.

OpenStudy (skullpatrol):